OpenStudy (anonymous):

the price of a bag is £12 in a sale; its normal price is £32. write the ratio of the sale price to the normal price in it's simplest form

OpenStudy (anonymous):

\[\frac{12}{32}=\frac{x}{100}\]

OpenStudy (anonymous):

\[12(100)=32(x)\]

OpenStudy (anonymous):

1200=32x

OpenStudy (anonymous):

1200/32=37.5=x

OpenStudy (anonymous):

it costs 37.5% of original cost

OpenStudy (anonymous):

wait u want ratio :P

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

wouldnt it be 8:3?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

no, wait, 3:8

OpenStudy (anonymous):

sale price:normal price

OpenStudy (anonymous):

12/32 = 3/8

OpenStudy (anonymous):

yeah